OpenStudy (anonymous):

write a possible explicit rule for the nth term of each sequence 1.5 , 3 , 4.5 , 6, 7.5

OpenStudy (reemii):

you have to find out if the sequence is rather an arithmetic sequence or a geometric sequence ("do I add something at every step?" VS "do I multiply by some number at every step?") what do you think here?

OpenStudy (anonymous):

add 1.5 every step

hartnn (hartnn):

thats correct! so this is an arithmetic sequence since we add 1.5 every step, d = common difference = 1.5 a1 = 1st term = 1.5 just use this formula \(\Large a_n = a_1 + (n-1)d\)

OpenStudy (anonymous):

an= 1.5 + (n-1)1.5

hartnn (hartnn):

thats correct :) now simplify it

hartnn (hartnn):

not actually.... distribute 1.5 to (n-1) \(1.5(n-1) = 1.5 \times n - 1.5 \times 1\) got that ??

hartnn (hartnn):

now try to simplify again

OpenStudy (anonymous):

i dont know , im hopeless ;o

hartnn (hartnn):

\(a_n = 1.5 +1.5n -1.5\) look what gets cancelled!

OpenStudy (reemii):

even though it's not really asked (" write a possible explicit formula .."), it's nice sometimes to see that it can be written in a very simple form. \( 1.5 + (n-1)\times 1.5 = 1.5 + n\times1.5 - 1.5 = n\times 1.5\).

OpenStudy (anonymous):

you call that very simple ? i call that Spanish.

hartnn (hartnn):

did you notice 1.5 -1.5 becoming 0 ?

OpenStudy (anonymous):

so its an = o + 1.5n

OpenStudy (reemii):

correct

hartnn (hartnn):

which is just 1.5n !

hartnn (hartnn):

so, \(a_n = 1.5n\)

hartnn (hartnn):

thats it! :)

OpenStudy (reemii):

however if you stick with 1.5+ (n-1)1.5, it's just as correct.

OpenStudy (anonymous):

thanks , even tho you guys probably got irritated ... sorry though.!! <3

hartnn (hartnn):

no problem, you're learning! welcome ^_^